# Biharmonic nonlinear vector field equations in dimension 44

### v2

@@ 4 Some lemmas applicable to the solutions of (2)
Here we gather some lemmas applicable to the solutions of (2). We first recall a regularity result for the nonhomogeneous biharmonic equation that can be derived from the $L^{p}$ estimates in [18, (2.6)].
@@
Next, we state Pohozaev identity for solutions to (2) belonging to $\mathcal{C}$. This identity was established in [16, Theorem 1.2] for dimensions $d\geq 5$. Its proof is identical in dimension $4$.
@@
Let $v\in\mathcal{C}$ be a solution to (2) in $\mathcal{D}^{\prime}$. Then, $\int_{\mathbb{R}^{4}}G(v)=0$.
@@
We notice that since $v$ as well as $g(v)$ are bounded in $\mathbb{R}^{4}$, it follows from Lemma 9 that $v\in W^{4,q}_{loc}(\mathbb{R}^{4})$, $\forall q\geq 1$. Then, we reproduce the arguments in the proof of [16, Theorem 1.2]. ∎
@@
where $B_{r}=\{x\in\mathbb{R}^{4}:|x|<1\}$, and $\nu=(\nu_{1},\nu_{2},\nu_{3},\nu_{4})$ is the outward unit normal vector at $\partial B_{r}$. Moreover, since $|\nabla\phi||D^{2}\phi|\in L^{\frac{4}{3}}(\mathbb{R}^{4})$, there exists a sequence $r_{k}\to\infty$, such that $\lim_{k\to\infty}r_{k}\int_{\partial B_{r_{k}}}(|\nabla\phi||D^{2}\phi|)^{\frac{4}{3}}=0$. Therefore, we conclude that
$\Big|\int_{\partial B_{r_{k}}}(\nu_{j}\partial_{i}\phi\cdot\partial^{2}_{ij}\phi-\nu_{i}\partial_{i}\phi\cdot\partial^{2}_{jj}\phi)\Big|\leq\Big[\int_{\partial B_{r_{k}}}(|\nabla\phi||D^{2}\phi|)^{\frac{4}{3}}\Big]^{\frac{3}{4}}O(r^{\frac{3}{4}})\to 0,$
@@
On the one hand, it is clear that $\lim_{r\to\infty}\int_{B_{r}}\phi\cdot\Delta\phi=\int_{\mathbb{R}^{4}}\phi\cdot\Delta\phi$. On the other hand, we notice that since $|\nabla\phi||\phi|\in L^{\frac{4}{3}}(\mathbb{R}^{4})$, there exists a sequence $r_{k}\to\infty$, such that $\lim_{k\to\infty}r_{k}\int_{\partial B_{r_{k}}}\frac{\partial\phi}{\partial\nu}\cdot\phi=0$. Therefore, we have
$\Big|\int_{\partial B_{r_{k}}}\frac{\partial\phi}{\partial\nu}\cdot\phi\Big|\leq\Big[\int_{\partial B_{r_{k}}}(|\nabla\phi||\phi|)^{\frac{4}{3}}\Big]^{\frac{3}{4}}O(r^{\frac{3}{4}})\to 0,$
@@ 5 Conclusion
The method of constrained minimization applies to prove the existence of a ground state solution for both systems (1) and (2), in space dimensions greater or equal than that of the limiting case (i.e. $d\geq 2$ for (1), and $d\geq 4$ for (2)). In contrast with the second order system (1), none asymptotic condition on the potential $G$ is required for the fourth order system (2), since for the latter, the relevant class $\cal C$ of functions is contained in $L^{\infty}$. We also point out that if the space dimension is smaller than that of the limiting case, then the problem of finding nontrivial solutions converging asympototically to 0 is different. In space dimension $d=1$, such solutions of (1) are the homoclinic orbits (cf. for instance [2]). On the other hand, the construction of such solutions of (2) in space dimensions $d=1,2,3$, is to a large extent an open problem.

### v3

@@ 4 Some lemmas applicable to the solutions of (2)
Here, we gather some lemmas applicable to the solutions of (2). We first recall a regularity result for the nonhomogeneous biharmonic equation that can be derived from the $L^{p}$ estimates in [24, (2.6)].
@@
Next, we examine the regularity of solutions of (2) belonging to $\mathcal{C}$, and state the corresponding Pohozaev identity. This identity was established in [23, Theorem 1.2] for dimensions $d\geq 5$. Its proof is identical in dimension $4$.
@@
Assuming that assumptions (3)-(7) hold, let $v\in\mathcal{C}$ be a ditributional solution of (2). Then, we have
  • (i)
    $v\in C^{3,\alpha}_{loc}\cap W^{4,q}_{loc}$ for any $0<\alpha<1$ and $1\leq q<\infty$,
  • (ii)
    $\lim_{|x|\to\infty}v(x)=\lim_{|x|\to\infty}|\nabla v(x)|=0$,
  • (iii)
    $\int_{\mathbb{R}^{4}}G(v)=0$ (Pohozaev identity).
@@
(i) Let $v\in\mathcal{C}$ be a ditributional solution of (2). Since, $\mu(\{|v|>\epsilon/2\})<\infty$, we can see as in the proof of Lemma 1, that inequality (18) applies to $v$, for every $q>1$. In particular, given $q>1$, the norms $\|v\|_{L^{q}(B(x,1))}$ are uniformly bounded on unit balls ($\forall x\in\mathbb{R}^{4}$). Similarly, the norms $\|g(v)\|_{L^{q}(B(x,1))}$ are uniformly bounded on unit balls. To see this, we use the fact that $\int_{\mathbb{R}^{4}}A(k_{1}(|v|-\epsilon/2)^{+})<\infty$, for a constant $k_{1}>0$ (since $(|v|-\epsilon/2)^{+}\in W^{1,4}(\mathbb{R}^{4})$), together with the bounds $|g(v)|\leq k_{2}+A(\frac{k_{1}}{q}(|v|-\epsilon/2)^{+})$ and $|A(v)|^{q}\leq k_{3}+A(qv)$ holding in $\mathbb{R}^{m}$ for constants $k_{2},k_{3}>0$. Therefore, it follows from Lemma 9 that $v\in W^{4,q}_{loc}$, $\forall q>1$, and from the Sobolev embeddings that $v\in C^{3,\alpha}_{loc}$, $\forall\alpha\in(0,1)$.
(ii) Let $w=\Delta v\in L^{2}(\mathbb{R}^{4})$. Another consequence of what precedes is that the norms $\|w\|_{W^{2,2}(B(x,1/2))}$ are uniformly bounded ($\forall x\in\mathbb{R}^{4}$). This follows from the interior estimates of [17, Theorem 4.9] applied to $w$ and $\Delta w=g(v)$, since we have seen that the norms $\|g(v)\|_{L^{2}(B(x,1))}$ are uniformly bounded. Next, an application of [17, Theorem 4.11] to $v$ and $\Delta v=w$, shows that the norms $\|v\|_{W^{4,2}(B(x,1/4))}$ are uniformly bounded ($\forall x\in\mathbb{R}^{4}$), since the norms $\|v\|_{L^{2}(B(x,1))}$ are uniformly bounded. Therefore, we deduce from the Sobolev embeddings that $v$ as well as its first derivatives are bounded and uniformly continuous on $\mathbb{R}^{4}$.
To prove that $v$ converges asymptotically to 0, we first establish that the set $\{|v|>\epsilon\}$ is bounded. Indeed, if $|v(x_{n})|>\epsilon$ holds for a sequence $\{x_{n}\}\subset\mathbb{R}^{4}$ such that $\lim_{n\to\infty}|x_{n}|=\infty$, then in view of the uniform continuity of $v$, there exists $\delta>0$ such that $|v|>\epsilon/2$ holds on the balls $B(x_{n},\delta)$, which by passing to a subsequence can be assumed to be disjoint. Consequently, we obtain that $\mu(\{|v|>\epsilon/2\})=\infty$, a contradiction. Now, suppose that there exists a sequence $\{x_{n}\}\subset\mathbb{R}^{4}$ such that $\lim_{n\to\infty}|x_{n}|=\infty$, and $|v(x_{n})|>\eta$, for some $\eta\in(0,\epsilon)$. According to what precedes, we have $|v(x_{n})|\in[\eta,\epsilon]$ (for $n$ large enough). Then, in view of the uniform continuity of the function $\mathbb{R}^{4}\ni x\mapsto G(v(x))$ and (5), there exists $\delta>0$, such that $|G(v)|>\beta>0$ holds on the balls $B(x_{n},\delta)$, for some $\beta>0$. As previously, the balls $B(x_{n},\delta)$ can be assumed to be disjoint, thus we get $G(v)\notin L^{1}(\mathbb{R}^{4})$, a contradiction. Therefore, we proved that $\lim_{|x|\to\infty}v(x)=0$. In a similar way, one can see that $\lim_{|x|\to\infty}|\nabla v(x)|=0$.
(iii) To derive Pohozaev identity, reproduce the arguments in the proof of [23, Theorem 1.2]. ∎
@@
where $B_{r}=\{x\in\mathbb{R}^{4}:|x|<r\}$, and $\nu=(\nu_{1},\nu_{2},\nu_{3},\nu_{4})$ is the outward unit normal vector at $\partial B_{r}$. Moreover, since $|\nabla\phi||D^{2}\phi|\in L^{\frac{4}{3}}(\mathbb{R}^{4})$, there exists a sequence $r_{k}\to\infty$, such that $\lim_{k\to\infty}r_{k}\int_{\partial B_{r_{k}}}(|\nabla\phi||D^{2}\phi|)^{\frac{4}{3}}=0$. Therefore, we conclude that
$\Big|\int_{\partial B_{r_{k}}}(\nu_{j}\partial_{i}\phi\cdot\partial^{2}_{ij}\phi-\nu_{i}\partial_{i}\phi\cdot\partial^{2}_{jj}\phi)\Big|\leq\Big[\int_{\partial B_{r_{k}}}(|\nabla\phi||D^{2}\phi|)^{\frac{4}{3}}\Big]^{\frac{3}{4}}O(r_{k}^{\frac{3}{4}})\to 0,$
@@
On the one hand, it is clear that $\lim_{r\to\infty}\int_{B_{r}}\phi\cdot\Delta\phi=\int_{\mathbb{R}^{4}}\phi\cdot\Delta\phi$. On the other hand, we notice that since $|\nabla\phi||\phi|\in L^{\frac{4}{3}}(\mathbb{R}^{4})$, there exists a sequence $r_{k}\to\infty$, such that $\lim_{k\to\infty}r_{k}\int_{\partial B_{r_{k}}}(|\nabla\phi||\phi|)^{\frac{4}{3}}=0$. Therefore, we have
$\Big|\int_{\partial B_{r_{k}}}\frac{\partial\phi}{\partial\nu}\cdot\phi\Big|\leq\Big[\int_{\partial B_{r_{k}}}(|\nabla\phi||\phi|)^{\frac{4}{3}}\Big]^{\frac{3}{4}}O(r_{k}^{\frac{3}{4}})\to 0,$
@@
Remark 6.
It follows from Lemma 11 that given the potential $G(u)=|u|^{2}\ln|u|$, we have $\mathcal{C}=\{u\in H^{2}(\mathbb{R}^{4}):\,G^{-}(u)\in L^{1}(\mathbb{R}^{4})\}$. Indeed, if $u\in\mathcal{C}$, it is clear that $u\in L^{2}(\mathbb{R}^{4})$, since $G(u)\in L^{1}(\mathbb{R}^{4})$, thus Lemma 11 implies that $\nabla u\in L^{2}(\mathbb{R}^{4})$, and $u\in H^{2}(\mathbb{R}^{4})$. Conversely, if $u\in H^{2}(\mathbb{R}^{4})$, we have $u,\,\nabla u\in H^{1}(\mathbb{R}^{4})\subset L^{4}(\mathbb{R}^{4})$, and $\mu(\{|u|>\eta\})<\infty$, $\forall\eta>0$, since $u\in L^{2}(\mathbb{R}^{4})$. Moreover, it holds $G^{+}(u)\in L^{1}(\mathbb{R}^{4})$, in view of the bound $G^{+}(u)\leq|u|^{4}$.
@@ 5 Conclusion
The method of constrained minimization applies to prove the existence of a ground state solution for both systems (1) and (2), in space dimensions greater or equal than that of the limiting case (i.e. $d\geq 2$ for (1), and $d\geq 4$ for (2)). In dimension $d=2$, the ground state solution of the scalar second order equation (1) ($m=1$) was constructed in [7], assuming that the potential $G$ satisfies an exponential growth condition. This hypothesis is required to obtain compactness properties from the Trudinger inequality. Similarly, for the fourth order system (2) in dimension $d=4$, an exponential growth condition is assumed for $G$. We also point out that if the space dimension is smaller than that of the limiting case, then the problem of finding nontrivial solutions converging asympototically to 0 is different. In space dimension $d=1$, such solutions of (1) are the homoclinic orbits (cf. for instance [5]). On the other hand, the construction of such solutions of (2) in space dimensions $d=1,2,3$, is to a large extent an open problem.